OpenStudy (anonymous):

The equation of a hyperbola is given: a. Identify the coordinates of the center of the hyperbola. b. Find the length of the transverse and conjugate axes. c. Find the slopes of the asymptotes. ( it would be easier to graph if it was in "Y=" form)

OpenStudy (anonymous):

\[\frac{ y^2 }{ 25 }-\frac{ (x-6)^2 }{ 144 }=1\]

OpenStudy (mertsj):

Do you know the center?

OpenStudy (mertsj):

\[\frac{(y-k)^2}{a^2}-\frac{(x-h)^2}{b^2}=1\]

OpenStudy (mertsj):

Center is (h,k)

OpenStudy (anonymous):

would it be (6,0) @Mertsj